Determine the Laplace transform of the given generalized function. 480t - 4) L{48(t - 4)}(s) =
Answer #1

dirac delta funn is denoted by & Lt-a) and is 817-4) = lim &lt-), there defined by Eto do Ct-a)= for astsate otherwise = 0 HE4+6 1 14 8lt-4)) (s) lim (12) E-TO E = 4 lim 1 -st 14+E 45 = 4 lim Eto & [ e-$(4+6) -S -S 24 lim -se SE [e - As e-SE = 4 lim